# On an entropy inequality for quadratic forms and applications

### v2

@@ 7 Examples
In this section, we show details of some examples. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
@@
Indeed, let $m=m(n)\in\mathbb{N}$ and $l=l(n)\in\mathbb{N}\cup\{0\}$ be parameters to be chosen later. We suppose that $(X,Y)$ has the uniform distribution on $A\times A$, with
@@
Denote by $\mathcal{L}^{d}$ the $d$-dimensional Lesbegue measure. Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have Note that $X$ has uniform distribution on $A$, i.e.,
@@
for any Borel set $E$.
In the following, we will proof the following results.
@@
and $Y$ has uniform distribution on $0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is $(1/2,O(2^{m}))$-Frostman. We take $s=1/2$ and consider $n=2m$.
Claim. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=ns+O(1)$.
In this example, we do not obtain a positive constant $\epsilon$ in Theorem 1.2. The reason is that the Frostman constant for $Y$ grows with $m$. More concretely, given $C$ of magnitude $2^{m}$, when $n=2m$ is of magnitude $\log C$, the two entropies may be small simultaneously. Therefore, in the conclusion of Theorem 1.2, it is essential to require $n>N$, where $N$ depends on the Frostman constant $C$.
@@
The conclusion follows by noting that $H_{n}(X^{\prime})=m=ns$. ∎
@@
This example is an variant of Example 7.4. Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}^{2}$. Note that $X$ and $Y$ are dependent. The marginal distributions of $X$ and $Y$ are both the uniform distribution on $\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$. So both $X$ and $Y$ are $(1/2,O(1))$-Frostman. Let $W_{1}=(X_{1},Y_{1}),\,(X_{2},Y_{2})$ be i.i.d. copies of $(X,Y)$. We consider $n\leq 2m$.
Claim 1. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq m+O(1)$.
@@
If we take $n=2m$ in Claim 1, then $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq n/2+O(1)$. In such a situation, the two entropies can be small simultaneously. Therefore, the requirement of independence of $X$ and $Y$ is essential in Theorem 1.2.
Moreover, for the random variable $W$, its Frostman constant $C$ can grow with $m$. If we consider $n=2m$ in Theorem 2.30, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2. That is to say, the term $\log C$ in Theorem 2.30 is essential.
@@
$range(X^{\prime}+Y^{\prime})\subseteq\{j2^{-m+1}:\,j=0,1,\ldots,2^{m}-1\},\text{ and}$
$range(X^{\prime 2}+Y^{\prime 2})\subseteq\{j^{2}2^{-2m+1}:\,j=0,1,\ldots,2^{m}-1\}.$
@@
$\max\{H_{n}(X^{\prime}+Y^{\prime}),\,H_{n}(X^{\prime 2}+Y^{\prime 2})\}\leq m.$
@@
This example is a further variant of Examples 7.4 and 7.5. Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}\times\big{(}-j2^{-m},\,-j2^{-m}+2^{-2m}\big{)}$. Then the marginal distributions of $X$ and $Y$ are the uniform distribution on the sets $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$ and $A_{0}^{\prime}=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}-j2^{-m},\,-j2^{-m}+2^{-2m}\big{)}$, respectively. So both $X$ and $Y$ are $(1/2,O(1))$-Frostman.
Claim 1. The random variable $(X,Y)$ is jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(2))$-Frostman for some $(s_{1},s_{2})\neq(0,0)$.
Claim 2. Taking $n=2m$, one has $H_{n}(X+Y)=H_{n}(X^{2}-Y^{2})=O(1)$, and $H_{n}(X,Y)=m+O(1)$.
In particular, if we apply Theorem 1.4 with $(X,Y)$, take $\alpha=-\beta$, $s_{1}=s_{2}=s=1/4$ and $n=2m$, then the second and third entropies in the maximum-term nearly vanish. So the first term $(H_{n}(X)+H_{n}(Y))/2$ is essential. In this example, it is $m+O(1)$, which is obviously larger than $ns=m/2$. Furthermore, we can not substitute $\frac{1}{2}H(X,Y)=m/2$ for $(H_{n}(X)+H_{n}(Y))/2$ as in Theorem 1.3.
Proof of Claim 1.
Note that $\mathcal{L}^{2}(A)=2^{-3m}$ and $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be a Borel set and $B(y,r)$ be any ball. When $0<r<2^{-2m}$, we have
$P(X\in E,Y\in B(y,r))\ll\mathcal{L}^{1}(E\cap A_{0})\cdot r\cdot 2^{3m},$
while
$P(X\in E)\leq\mathcal{L}^{1}\cdot 2^{m}.$
It follows that
$P(X\in E,Y\in B(y,r))\ll r^{1/2}P(X\in E).$
For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, one has
and the other cases with $r_{2}<r_{1}$ can be obtained by symmetry. One can verify that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$.
For the conditional Frostman condition, we take $E=B(x,2^{-2m})$ and $F=B(y,2^{-2m})$ such that $E\times F\subseteq A$. Then
$P(X\in E)=2^{-m}=P(X\in E,\,Y\in B(x,2^{-2m}))=2^{-m}.$
The only possible positive $s_{2}$ satisfying $2^{-m}\leq 2^{-m}(2^{-2m})^{s_{2}}$ is 0. Now it is not hard to see that $(X,Y)$ is $(s_{1},O(1);s_{2},O(1))$-Frostman only when $(s_{1},s_{2})=(0,0)$. ∎
Proof of Claim 2.
We use similar arguments as in the proof of Claim 1 in Example 7.5. Let $(X^{\prime},Y^{\prime})=D_{2m}(X,Y)$. Then $(X^{\prime},Y^{\prime})$ is the discrete uniform distribution on $\{(j2^{-m},-j2^{-m}:\,j=0,1,\ldots,2^{m}-1\}$. So $H_{n}(X^{\prime},Y^{\prime})=m$. Moreover, $range(X^{\prime}+Y^{\prime})=range(X^{\prime 2}-Y^{\prime 2})=\{0\}$, which implies that $H_{n}(X^{\prime}+Y^{\prime})=H_{n}(X^{\prime 2}-Y^{\prime 2})=0$. Applying similar relation as in (66), one asserts that $H_{n}(X+Y)=H_{n}(X^{2}-Y^{2})=O(1)$. Similarly, $H_{n}(X,Y)=m+O(1)$. ∎
Example 7.7.
@@
$B=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},i2^{-m}+(i+1)2^{-2m}).$
Then, for each given $x\in[0,1)$ (or $y\in[0,1)$), the set $\{y:\,(x,y)\in B\}$ (or $\{x:\,(x,y)\in B\}$, respectively) is exactly an interval of length $2^{-2m}$. Now Let
$A=\bigcup\limits_{i^{\prime},j^{\prime}=0}^{2^{m}-1}(i2^{-m},j2^{-m})+2^{-2m}\cdot B.$
Here $2^{-2m}\cdot B=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in B\}$. Suppose that $(X,Y)$ is a random variable that has uniform distribution on $A$.
@@
One can check that both $X$ and $Y$ have the uniform distribution on the set $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right)$. By Example 7.1, they are $(1/2,O(1))$-Frostman. ∎
@@
By Lemma 2.21, we have the following statement: Let $k,l\in\mathbb{Q}$ be given. Then

### v3

@@ 5.2.2 Modifications to their argument
Proof of Theorem 1.6.
Similar arguments as above will be applied. We only mention the differences that is needed to make.
Since $\{y:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\gg\delta^{-s+\epsilon}$ for any $x\in A$, one has $G\gg\delta^{-2s+\epsilon}$. Also, one has $G\leq(\#A)^{2}\ll\delta^{-2s}$. As an alternative of (68), we have
$\delta^{2-2s+\epsilon}\ll_{\phi}\mathcal{L}^{2}(G_{\eta\delta})\ll_{\phi}\delta^{2-2s}.$
And we have
$\mathcal{E}_{\delta}(G_{\eta\delta})=\mathcal{E}_{\delta}(G)=\#G\leq(\#A)^{2}\ll\delta^{-2s}$
instead of (67).
Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}$. Let $E$ be any Borel set and $B(y,r)$ be any ball. We have
$P(X\in E)\gg_{\phi}\mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot\delta^{1-s+\epsilon}\cdot\delta^{2s-2}.$
And
$P(X\in E,Y\in B(y,r))\ll_{\phi}\begin{cases}\mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot r\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r\leq\delta,\\ \mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot Cr^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }\delta<r\leq 1,\\ \mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},&\text{if }r>1.\end{cases}$
In all the cases,
$P(X\in E,Y\in B(y,r))\ll C\delta^{-2\epsilon}r^{s}P(X\in E)$
So $Y$ is conditionally $(s,O_{\phi}(C\delta^{-2\epsilon}))$-Frostman given $X$. Similarly $X$ is also conditionally $(s,O_{\phi}(C\delta^{-2\epsilon}))$-Frostman given $Y$. Then $(X,Y)$ is conditionally $(s,O_{\phi}(C\delta^{-2\epsilon});s,O_{\phi}(C\delta^{-2\epsilon}))$-Frostman.
Now (69) can be replaced by
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$
for some $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)>0$, which results from Theorem 1.3’. Moreover, we substitute
$\log\#range_{+}(D_{n}(X,Y))\geq H_{n}(X,Y)$
and
$\mathcal{E}_{\delta}(G_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(G_{\eta\delta})\gg\#range_{+}(D_{n}(X,Y))$
for the first formula in (70) and (71), respectively. Hence (73) can be replaced by
$\max\left\{\mathcal{E}_{\delta}(G)^{1/2},\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}.$
Finally, since $\mathcal{E}_{\delta}(G)\ll\delta^{-2s}$, we conclude that
$\max\left\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$
whenever $\delta$ is sufficiently small depending on $\phi$, $s$ and $C$. ∎
Proof of Theorem 1.7.
We use similar arguments as in the proof of Theorem 1.5, and work with $G=B\times B$. Since $B\geq\delta^{\epsilon}\#A\geq\delta^{-s+\epsilon}$, one has $\mathcal{L}^{1}(B_{\eta\delta})\gg\delta^{1-s+\epsilon}$. Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}=B_{\eta\delta}\times B_{\eta\delta}$. It is not hard to see that $X$ and $Y$ are i.i.d.. Combining the non-concentration condition, one deduces that
$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\cdot\delta^{s-1-\epsilon},\quad&\text{if }0<r\leq\delta,\\ Cr^{s}\delta^{1-s}\cdot\delta^{s-1-\epsilon},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1,\end{cases}$
for any ball $B(x,r)$. In all the cases, the right-hand side can be bounded by $C\delta^{-\epsilon}r^{s}$. So $X$ is $(s,O_{\phi}(C\delta^{-\epsilon}))$-Frostman.
As a replacement of (69), Theorem 1.4’ shows that
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\epsilon^{\prime})+O_{s}(\log C)+O_{\phi,s}(1)$
for some $\epsilon^{\prime}=\epsilon^{\prime}(s)>0$ and for all $n\in\mathbb{N}$. And the similar conclusion as in (73) appears that
$\max\{\mathcal{E}_{\delta}(B+B),\,\mathcal{E}_{\delta}(\phi(B,B))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
for all $\delta>0$. ∎
Remark 6.1.
In the conclusions of Theorems 1.5-1.7, the term $C^{-O_{s}(1)}$ can be replaced by $C^{-O(1)}$ if either of the following two situations happens: (1) we are only considering the case $s\gg 1$; (2) the condition $A\subseteq[0,1]$ is replaced by $A\subseteq[1,2]$.
@@ 7 Examples
In this section, we show details of some examples. Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Here measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
@@
Indeed, let $m=m(n)\in\mathbb{N}$ and $l=l(n)\in\mathbb{N}\cup\{0\}$ be parameters to be chosen later. We suppose that $(X,Y)$ has the uniform distribution on $G=A\times A$, with
@@
Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have Note that $X$ has uniform distribution on $A$, i.e.,
@@
for any Borel set $E$. In the following, we will proof several results. The $o(1)$-terms are in the situation that $n\rightarrow\infty$.
@@
and $Y$ has uniform distribution on $0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is both $(1/2,O(2^{m}))$-Frostman and $(0,1)$-Frostman. We take $n=2m$.
Claim. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=n/2+O(1)$.
In this example, we do not obtain a positive constant $\epsilon$ in Theorem 1.4 if $(X,Y)$ is regarded as a marginally $(1/2,O(1);1/2,O(2^{m}))$-Frostman random variable, i.e., $s=1/2$. The reason is that the Frostman constant for $Y$ grows with $m$. More concretely, given $C$ of magnitude $2^{m}$, when $n=2m$ is of magnitude $\log C$, the two entropies may be small simultaneously. Therefore, in the conclusion of Theorem 1.4, it is essential to require $n>N$, where $N$ depends on the Frostman constant $C$. More concretely, the term $O_{s}(\log_{+}C_{1}C_{2})$ in the conclusion in Theorem 1.4’ is essential.
On the other hand, if we regard $(X,Y)$ as a marginally $(1/2,O(1);0,1)$-Frostman random variable, then $s=1/4$ and Theorem 1.4 works.
@@
The conclusion follows by noting that $H_{n}(X^{\prime})=m=n/2$. ∎
@@
This example is an variant of Example 7.4. Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $G=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}^{2}$. Note that $X$ and $Y$ are dependent. Let $W_{1}=(X_{1},Y_{1}),\,(X_{2},Y_{2})$ be i.i.d. copies of $(X,Y)$. We consider $m\leq n\leq 2m$.
Claim 1. The random variable $(X,Y)$ is marginally marginally $(1/2,O(1);1/2,O(1))$-Frostman, and jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(1))$-Frostman for any $(s_{1},s_{2})\neq(0,0)$.
Claim 2. We have $\max\{H_{2m}(X+Y),\,H_{2m}(X^{2}+Y^{2})\}\leq m+O(1)$, $H_{2m}(X,Y)=m+O(1)$ and $H_{2m}(X-Y)=H_{2m}(X^{2}-Y^{2})=O(1)$.
@@
Now we consider the situation $n=2m$. By Claim 2, one has $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq n/2+O(1)$. In such a situation, the two entropies can be small simultaneously. Therefore, the requirement of independence of $X$ and $Y$ in Theorem 1.4 is essential.
Moreover, if we apply Theorem 1.2 with the random variable (X,Y)(X,-Y), and take $\alpha=1$, $\beta=-1$, $s_{1}=s_{2}=s=1/4$, then
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X-Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon)$
for some $\epsilon>0$. By Claim 2, the second and third entropies in above maximum-term nearly vanish. So the first term $(H_{n}(X)+H_{n}(Y))/2$ is essential. In this example, it is $m+O(1)$, which is obviously larger than $ns=m/2$. Furthermore, we can not substitute $\frac{1}{2}H(X,Y)=m/2$ for $(H_{n}(X)+H_{n}(Y))/2$ as in Theorem 1.3.
Furthermore, for the random variable $W$, its Frostman constant $C$ can grow with $m$. If we consider $n=2m$ in Theorem 2.38, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2. That is to say, the term $\log_{+}C$ in Theorem 2.38 is essential.
Proof of Claim 1.
The marginal distributions of $X$ and $Y$ are both the uniform distribution on $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$. Example 7.1 shows that $(X,Y)$ is marginally $(1/2,O(1);1/2,O(1))$-Frostman.
Note that $\mathcal{L}^{2}(G)=2^{-3m}$ and $\mathcal{L}^{1}(A)=2^{-m}$. For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, one has
and the other cases with $r_{2}<r_{1}$ can be obtained by symmetry. One can verify that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$. So $(X,Y)$ is jointly $(1/4,1/4,O(1))$-Frostman.
For the conditional Frostman condition, one can check that
$P(X\in[0,2^{-2m}))=P(X\in[0,2^{-2m}),\,Y\in B(0,2^{-2m}))=2^{-m}.$
The only possible positive $s_{2}$ satisfying $2^{-m}\ll 2^{-m}(2^{-2m})^{s_{2}}$ for all sufficiently large $m$ is 0. Now it is not hard to see that $(X,Y)$ is $(s_{1},o(1);s_{2},O(1))$-Frostman only when $(s_{1},s_{2})=(0,0)$. ∎
@@
$range_{+}(X^{\prime}+Y^{\prime})\subseteq\{j2^{-m+1}:\,j=0,1,\ldots,2^{m}-1\},$
$range_{+}(X^{\prime 2}+Y^{\prime 2})\subseteq\{j^{2}2^{-2m+1}:\,j=0,1,\ldots,2^{m}-1\},$
and
$range_{+}(X^{\prime}-Y^{\prime})=range_{+}(X^{\prime 2}-Y^{\prime 2})=\{0\}.$
@@
$\max\{H_{n}(X^{\prime}+Y^{\prime}),\,H_{n}(X^{\prime 2}+Y^{\prime 2})\}\leq m,$
and
$H_{n}(X^{\prime}-Y^{\prime})=H_{n}(X^{\prime 2}-Y^{\prime 2})=0.$
The conclusion then follows. ∎
@@
$G^{\prime}=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},i2^{-m}+(i+1)2^{-2m}).$
Then, for each given $x\in[0,1)$ (or $y\in[0,1)$), the set $\{y:\,(x,y)\in G^{\prime}\}$ (or $\{x:\,(x,y)\in G^{\prime}\}$, respectively) is exactly an interval of length $2^{-2m}$. Now Let
$G=\bigcup\limits_{i^{\prime},j^{\prime}=0}^{2^{m}-1}(i2^{-m},j2^{-m})+2^{-2m}\cdot G^{\prime}.$
Here $2^{-2m}\cdot G^{\prime}=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in G^{\prime}\}$. Suppose that $(X,Y)$ is a random variable that has uniform distribution on $G$.
@@
One can check that both $X$ and $Y$ have the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right)$. By Example 7.1, they are $(1/2,O(1))$-Frostman. ∎
@@
By Theorem 2.23, we have the following statement: Let $k,l\in\mathbb{Q}$ be given. Then